What quadratic equation would have solutions x= 2/3 and x=1/5
Answer by ankor@dixie-net.com(22740)

What quadratic equation would have solutions x= 2/3 and x=1/5
x = 2/3; from the factor: (3x - 2)
and
x = 1/5; from the factor: (5x - 1)
:
FOIL
(3x - 2)(5x-1) = 15x^2 - 3x - 10x + 2 = (15x^2 - 13x + 2) is the equation